OpenStudy (anonymous):

use the quadratic formula to solve the equation 3x^2-7x-5=0

OpenStudy (bahrom7893):

okay here's the quadratic: first find the discriminant: D = b^2-4ac

OpenStudy (bahrom7893):

your equation is in the form: ax^2+bx+c=0 where: a=3, b=-7,c=-5 D = b^2-4ac = (-7)^2-4*(3)*(-5) = 49 + 60 = 109

OpenStudy (bahrom7893):

x1 = ( -b-sqrt(D) ) / (2a) = ( -(-7)- sqrt(109) )/(2*3)

OpenStudy (bahrom7893):

x1 = (7-sqrt(109))/6 x2 is ( -b+sqrt(D) ) / (2a), so: x2 = (7+sqrt(109))/6

OpenStudy (bahrom7893):

just use a calculator to estimate those..

OpenStudy (anonymous):

I still need a little help..

OpenStudy (bahrom7893):

with what?

OpenStudy (anonymous):

I do not have the proper tools to find the answer or else I have no clue what is going on in this problem

OpenStudy (anonymous):

to solve \[ax^2+bx+c=0\] use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

The formula for solutions of a quadratic ax^2 + bx +c = 0 is \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Plug the numbers into the formula and you get bahrom's result.